OpenStudy (anonymous):

find an equation of the line containing the given pair of points (-2,-1) (-6,-7)

OpenStudy (anonymous):

m=y2-y1/x2-x1 -7-(-1)/-6-(-2) -6/-4 m=1.5 y=mx+b -1=1.5(-2)+b b=2 y=1.5x+2

OpenStudy (anonymous):

y=kx+b through(-2.-1) (-6,-7) so {-1=-2k+b {-7=-6k+b so 4k=6 k=6/4=3/2 so b=2 s0 Y=(3/2)x+2

OpenStudy (anonymous):

lol good answers